OpenStudy (anonymous):

PLEASE EXPLAIN AND HELP? A garden has width sqrt13 and length 7sqrt13. What is the perimeter of the garden in simplest radical form?

OpenStudy (anonymous):

The perimeter is simply all of the sides added together. It helps to draw a picture:  \[\sqrt{13}+7\sqrt{13}+\sqrt{13}+7\sqrt{13}\] Since they all have the same radicand, you can add them all together. \[16\sqrt{13}\] This is the final awnser as you cannot simply further.